# Catastrophic Goodhart: regularizing RLHF with KL divergence does not mitigate heavy-tailed reward misspecification

### v1

@@ Abstract
When applying reinforcement learning from human feedback (RLHF), the reward is learned from data and, therefore, always has some error. It is common to mitigate this by regularizing the policy with KL divergence from a base model, with the hope that balancing reward with regularization will achieve desirable outcomes despite this reward misspecification. We show that when the reward function has light-tailed error, optimal policies under less restrictive KL penalties achieve arbitrarily high utility. However, if error is heavy-tailed, some policies obtain arbitrarily high reward despite achieving no more utility than the base model—a phenomenon we call catastrophic Goodhart. We adapt a discrete optimization method to measure the tails of reward models, finding that they are consistent with light-tailed error. However, the pervasiveness of heavy-tailed distributions in many real-world applications indicates that future sources of RL reward could have heavy-tailed error, increasing the likelihood of reward hacking even with KL regularization.
Machine Learning, ICML
@@ 1 Introduction
Kullback-Leibler (KL) divergence constraints in reinforcement learning (RL) are employed to stay in regimes where the objective is accurate enough. Some on-policy (Schulman et al., 2015, 2017) and many off-policy (Abdolmaleki et al., 2018; Jaques et al., 2019) policy gradient algorithms employ KL constraints or penalties during optimization to prevent the policy from deviating too much from the data collection distribution. This ensures that estimates of each action’s advantage are reliable enough to update the policy in a helpful way.
Reinforcement learning from human feedback (Christiano et al., 2017; Ziegler et al., 2020, RLHF) is a very popular method to induce desirable behavior in language models. RLHF starts with a base pre-trained model, then learns a reward function from human annotator data. Next, it trains an RL policy to maximize this reward, while penalizing high KL divergence from the policy to the base model. RLHF uses an on-policy algorithm and has accurate advantages, but the reward function is always somewhat misspecified compared to desired behavior, due to insufficient data, human biases, and other factors.
The main purpose of the KL penalty in RLHF is to limit the consequence of reward modeling errors by keeping the policy within a distribution similar to that on which it was trained. Ideally, in the low-KL regime the reward model’s errors are small enough that it provides correct updates to the base model. Gao et al. (2023) empirically supports this view: if the KL divergence in RLHF is allowed to grow too much, with a misspecified reward, the model’s performance on the true utility starts to decrease.
@@
Empirically, open-source language reward models seem to be light-tailed, which does not imply light-tailed errors but suggests it (Section 5). However, the errors are likely not independent and, given the prevalence of heavy-tailed distributions in the real world, error in future reward models may also be heavy-tailed. In any case, the present success of RLHF with misspecified rewards cannot be explained solely by the KL regularization in its objective.
@@ 2.1 KL divergence and KL regularization
Recall that KL divergence between two distributions P and Q is defined as $D_{\mathrm{KL}}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right)$.
@@
In RLHF, it is common to use the regularization term $\beta D_{KL}\left(\pi\|\pi_{0}\right)$ to prevent the learned policy from deviating too much from the base policy, which can prevent unstable behavior or overfitting to the reward model. If our reward model gives reward $U$, then the optimal policy for RLHF with a KL penalty is $\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}\left(\pi\|\pi_{0}\right).$
Often the regularization parameter $\beta$ is dynamically adjusted to keep the $D_{KL}$ near some target value (Ziegler et al., 2020).
@@ 2.2 Heavy-tailed distributions
A distribution $P$ over $\mathbb{R}$ with cumulative distribution function (CDF) $F_{P}$ is heavy-tailed if its tail function $\bar{F}_{P}(x)\triangleq 1-F_{P}(x)$ satisfies $\lim_{x\rightarrow\infty}e^{tx}\bar{F}(x)=\infty$ for all $t>0$. Heavy-tailed distributions are well-known in statistics to have a higher probability of producing a single extreme value. For example, if the sum of two independent variables from heavy-tailed distributions is large, it is most likely due to one extreme sample rather than two equally large samples. (Wierman, 2013)
@@ 2.3 Reward misspecification and Goodhart’s Law
Reward misspecification has caused low-utility outcomes in practice; for example, in (Clark & Amodei, 2016), an RL agent trained to play a racing videogame according to a misspecified reward function achieves a high score while failing to complete the course.
Gao et al. (2023) introduce the concept of “overoptimization”: optimizing for a proxy objective decreases performance according to the true objective. This raises the question: in general, when RLHF reward is misspecified, when does the optimal policy produce high utility?
@@
Reward misspecification is also studied by (Lambert & Calandra, 2024), (Laidlaw et al., 2024), and others. Laidlaw et al show that a KL penalty between action distributions can be ineffective, and propose instead regularizing state occupancy measure. Our results show an inherent weakness of KL divergence, including when applied to state occupancy measure.
@@ 3 Theoretical results
When applying KL regularization, the trained model is regularized towards some base policy $\pi_{0}$. One would hope that a KL penalty can produce good outcomes even in the case of reward misspecification; that is, if the reward $U$ is the sum of true utility $V$ and an error term $X$, we would hope that optimal policies under a KL penalty achieve high $V$ even if the magnitude of $X$ is large. We show that this is not always the case: Corollary 1 of Theorems 1, 2, and 3 establishes that when $X(\pi_{0})$ is heavy-tailed, there are arbitrarily well-performing policies $\pi$ with $\mathbb{E}_{\pi}[V]\approx\mathbb{E}_{\pi_{0}}[V]$. However, Theorem 4 shows that when error is light-tailed and independent of $V$, the optimal policy under a KL penalty results in $V>0$, and $V$ can be made arbitrarily large. Thus, the tails of the error distribution are crucial in determining how much utility will result from optimization towards an imperfect proxy.
Theorems 5 and 8 (Section B of the appendix) show that the relationship of catastrophic Goodhart to heavy-tailed error is not just a quirk of KL divergence by using a different model of optimization based on conditioning on high reward values. Under this model (and given additional regularity conditions), it is also true that heavy-tailed error results in catastrophic Goodhart, and light-tailed error plus independence results in arbitrarily large utility. All proofs are in the appendix.
3.1 Heavy-tailed distributions
@@ 3.2 RLHF with KL penalty under heavy-tailed return distribution
  • There is a set of sink states $E\subseteq\mathcal{S}$ that terminate a trajectory, which is disjoint from the set of start states.
@@
3.3 If V is light-tailed, $\mathbb{E}_{P}[V]\!-\!E_{Q}[V]\to 0$ as $D_{KL}\!\to\!0$
Theorem 3.
If $V$ is light-tailed and $d=D_{KL}(P\|Q)$ is bounded, then $\mathbb{E}_{P}[V]$ is bounded, and $\mathbb{E}_{P}[V]-\mathbb{E}_{Q}[V]\to 0$ as $d\to 0$.
@@ 3.4 Light-tailed + independence imply $\mathbb{E}[V]\to\infty$
If $U=X+V$ with $X$ and $V$ both light-tailed, and the distribution of U is continuous, and $\pi^{*}(\beta)\triangleq\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}(\pi,\pi_{0})$, then $\lim_{\beta\to 0^{+}}\mathbb{E}[V(\pi^{*}(\beta))]=\infty$.
4 Experimental Methodology
@@
Our theoretical results now raise the question of whether the error in reward models is heavy-tailed or light-tailed in practice.  [^1] Note that distributions over a finite set are bounded and cannot be heavy-tailed in a technical sense, and models with a finite context window have a finite input space. We say that a distribution of reward or error is heavy-tailed if it is well-modeled by a heavy-tailed distribution on its support. If we observe the reward distribution to be light-tailed, this is a strong indication that error is light-tailed.  [^2] It is possible for $U$ to be light-tailed while $X$ and $V$ are both heavy-tailed, but this is unusual and we do not expect it to happen in practice.
@@
To empirically test whether the reward is heavy-tailed, we consider two lines of evidence: examining the distributions directly through random sampling and temperature-1 sampling, and finding adversarial token sequences that get high rewards. We examine one small and one medium reward model that performed reasonably well on RewardBench (Lambert et al., 2023). The small model is an OpenAssistant model based on Pythia 1.4B, and the medium model is Starling 7B-alpha (Zhu et al., 2023) [^3] https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha.
@@
For random sampling, we sample 30000 length-1024 sequences of uniformly random tokens and observe the distribution of rewards assigned by both Pythia 1.4B and Llama 7B-chat. We also use Llama 7B-chat to generate 16000 length-133 sequences and observe the distribution of rewards assigned by Starling 7B-alpha.
@@
Because sampling is inefficient at probing the extreme tail, we also find token sequences that optimize Starling 7B-alpha for reward. We considered Greedy Coordinate Gradient (GCG) from (Zou et al., 2023), a method used to find adversarial suffixes that circumvent jailbreaking, but decided on a faster version of GCG called Accelerated Coordinate Gradient (ACG) from (Haize Labs, 2024). See Table 1 for ACG hyperparameters.
@@
5 Results
When sampling token sequences, both the Pythia model on random inputs (Figure 5) and Starling 7B-alpha on Llama-generated inputs (Figure 2) appear approximately normal and, therefore, light-tailed. Starling on random inputs (Figure 1 is ambiguous, with the exponential Q-Q plot having an outlier that could indicate a heavy-tailed distribution, but the Hill estimator is consistent with a light-tailed distribution. Because Llama-7B-chat is a more reasonable base model than a completely random policy, we believe that Starling 7B-alpha is more likely to be light-tailed for the purposes of our theoretical results.
@@ 5 Results
When $\alpha$ is small but much larger than $q(x)$, we approximate this to first order as $D_{KL}(P\|Q)\approx\alpha\log\left(\frac{\alpha}{q(x)}\right)$. In Theorems 1 and 2, we prove that when the error is sufficiently heavy-tailed, a policy that gets extremely large reward a small fraction of the time will achieve high expected reward with low KL divergence. This is not the case here because the rewards achieved through ACG were small and the log-probabilities extremely negative. For example, a policy that matches Llama 2-chat’s base reward 99% of the time and uses the highest-reward input generated by ACG $\alpha=$1% of the time will have KL divergence from Llama 2-chat of $\alpha(\log(\alpha)-1339.70)=13.35$ nats, but reward only about $\alpha*(2.2377-0.3329)=0.02571$ greater than the base model, far less than can be obtained with the same KL divergence by conditioning.
@@ 6.1 How likely is catastrophic Goodhart?
  • If we cap KL to a finite value (or dynamically adjust the KL penalty to target a finite KL, as done in Ziegler et al. (2020), then $\mathbb{E}[V]$ is also upper bounded by a finite value (see Theorem 3), and we think it is likely that $\mathbb{E}[V]\approx 0$. Consider a toy model where an AI can adjust three parameters: true quality $V$ of responses, frequency of reward hacking (producing actions with extremely high X), and severity of hacking (value of X on those actions). Adjusting the policy to increase $\mathbb{E}[U]$ without increasing KL increase the severity of hacking while decreasing either frequency of hacking or quality of responses. When $E[U]$ is already large, decreasing quality has much better returns than decreasing frequency. This is similar to Theorems 5, 8 about hard-threshold optimization.
  • Any way we maximize $\mathbb{E}[U(\pi)]-\beta D_{KL}\left(\pi,\pi_{0}\right)$ results in very large values of $\mathbb{E}[U(\pi)]$, and there are a number of arguments that extreme optimization for an imperfect proxy can result in decreased utility due to tradeoffs between $X$ and $V$; e.g., the constrained resource scenario in (Zhuang & Hadfield-Menell, 2021).

### v2

@@ Abstract
When applying reinforcement learning from human feedback (RLHF), the reward is learned from data and, therefore, always has some error. It is common to mitigate this by regularizing the policy with KL divergence from a base model, with the hope that balancing reward with regularization will achieve desirable outcomes despite this reward misspecification. We show that when the reward function has light-tailed error, optimal policies under less restrictive KL penalties achieve arbitrarily high utility. However, if error is heavy-tailed, some policies obtain arbitrarily high reward despite achieving no more utility than the base model–a phenomenon we call catastrophic Goodhart. We adapt a discrete optimization method to measure the tails of reward models, finding that they are consistent with light-tailed error. However, the pervasiveness of heavy-tailed distributions in many real-world applications indicates that future sources of RL reward could have heavy-tailed error, increasing the likelihood of reward hacking even with KL regularization.
@@ 1 Introduction
Kullback-Leibler (KL) divergence constraints in reinforcement learning (RL) are employed to stay in regimes where the objective is accurate enough. Some on-policy (Schulman et al., , 2015, 2017) and many off-policy (Abdolmaleki et al., , 2018; Jaques et al., , 2019) policy gradient algorithms employ KL constraints or penalties during optimization to prevent the policy from deviating too much from the data collection distribution. This ensures that estimates of each action’s advantage are reliable enough to update the policy in a helpful way.
Reinforcement learning from human feedback (Christiano et al., , 2017; Ziegler et al., , 2020, RLHF) is a very popular method to induce desirable behavior in language models. RLHF starts with a base pre-trained model, then learns a reward function from human annotator data. Next, it trains an RL policy to maximize this reward, while penalizing high KL divergence from the policy to the base model. RLHF uses an on-policy algorithm and has accurate advantages, but the reward function is always somewhat misspecified compared to desired behavior, due to insufficient data, human biases, and other factors.
The main purpose of the KL penalty in RLHF is to limit the consequence of reward modeling errors by keeping the policy within a distribution similar to that on which it was trained. Ideally, in the low-KL regime the reward model’s errors are small enough that it provides correct updates to the base model. Gao et al., (2023) empirically supports this view: if the KL divergence in RLHF is allowed to grow too much, with a misspecified reward, the model’s performance on the true utility starts to decrease.
@@
Empirically, open-source language reward models seem to be light-tailed, which does not imply light-tailed errors but suggests it (Section 4.1). However, the errors are likely not independent and, given the prevalence of heavy-tailed distributions in the real world, error in future reward models may also be heavy-tailed. In any case, the present success of RLHF with misspecified rewards cannot be explained solely by the KL regularization in its objective.
@@ 2.1 KL divergence and KL regularization
Recall that KL divergence between two distributions P and Q is defined as
$D_{\mathrm{KL}}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\left(\frac{P(x)}{Q(x)}\right).$
@@
In RLHF, it is common to use the regularization term $\beta D_{KL}\left(\pi\|\pi_{0}\right)$ to prevent the learned policy from deviating too much from the base policy, which can prevent unstable behavior or overfitting to the reward model. If our reward model gives reward $U$, then the optimal policy for RLHF with a KL penalty is
$\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}\left(\pi\|\pi_{0}\right).$
Often the regularization parameter $\beta$ is dynamically adjusted to keep the $D_{KL}$ near some target value (Ziegler et al., , 2020).
@@ 2.2 Heavy-tailed distributions
A distribution $P$ over $\mathbb{R}$ with cumulative distribution function (CDF) $F_{P}$ is heavy-tailed if its tail function $\bar{F}_{P}(x)\triangleq 1-F_{P}(x)$ satisfies
$\lim_{x\rightarrow\infty}e^{tx}\bar{F}(x)=\infty\qquad\text{for all}t>0.$
Heavy-tailed distributions are well-known in statistics to have a higher probability of producing a single extreme value. For example, if the sum of two independent variables from heavy-tailed distributions is large, it is most likely due to one extreme sample rather than two equally large samples. (Wierman, , 2013)
@@ 2.3 Reward misspecification and Goodhart’s Law
Reward misspecification has caused low-utility outcomes in practice; for example, in (Clark and Amodei, , 2016), an RL agent trained to play a racing videogame according to a misspecified reward function achieves a high score while failing to complete the course.
Gao et al., (2023) introduce the concept of “overoptimization”: optimizing for a proxy objective decreases performance according to the true objective. This raises the question: in general, when RLHF reward is misspecified, when does the optimal policy produce high utility?
@@
Reward misspecification is also studied by (Lambert and Calandra, , 2024), (Laidlaw et al., , 2024), and others. Laidlaw et al show that a KL penalty between action distributions can be ineffective, and propose instead regularizing state occupancy measure. Our results show an inherent weakness of KL divergence, including when applied to state occupancy measure.
We prove that in many cases, $V\to 0$ in the limit of optimization for some proxy $U$. We call this phenomenon “catastrophic Goodhart”, after Goodhart’s law: “when a measure becomes a target, it ceases to be a good measure” (Strathern, , 1997). In these cases, the end result of optimizing for a proxy of $V$ is no better than not optimizing at all. However, in other cases, $V\to\infty$ despite some reward misspecification; in these cases the reward misspecification is not severe enough to prevent good outcomes.
@@ 3 Theoretical results
When applying KL regularization, the trained model is regularized towards some base policy $\pi_{0}$. One would hope that a KL penalty can produce good outcomes even in the case of reward misspecification; that is, if the reward $U$ is the sum of true utility $V$ and an error term $X$, we would hope that optimal policies under a KL penalty achieve high $V$ even if the magnitude of $X$ is large. We show that this is not always the case: Corollary 1 of Theorems 1, 3, and 2 establishes that when $X(\pi_{0})$ is heavy-tailed, there are arbitrarily well-performing policies $\pi$ with $\mathbb{E}_{\pi}[V]\approx\mathbb{E}_{\pi_{0}}[V]$. However, Theorem 4 shows that when error is light-tailed and independent of $V$, the optimal policy under a KL penalty results in $V>0$, and $V$ can be made arbitrarily large. Thus, the tails of the error distribution are crucial in determining how much utility will result from optimization towards an imperfect proxy.
Theorems 5 and 6 (Section 3.4) show that the relationship of catastrophic Goodhart to heavy-tailed error is not just a quirk of KL divergence by using a different model of optimization based on conditioning on high reward values. Under this model (and given additional regularity conditions), it is also true that heavy-tailed error results in catastrophic Goodhart and light-tailed error plus independence results in arbitrarily large utility. All proofs are in the appendix.
3.1 KL divergence on heavy- and light-tailed distributions
@@
Intuitively, in a heavy-tailed distribution, events with extremely high $x$ are not very rare, so you don’t pay much of a KL penalty to upweight them so they happen about $1/x$ of the time.
Theorem 2.
However, if the distribution $Q$ is light-tailed and $d=D_{KL}(P\|Q)$ is bounded, then $\mathbb{\mu}_{P}$ is bounded, and $\mu_{P}-\mu_{Q}\to 0$ as $d\to 0$.
@@ 3.2 RLHF with KL penalty under heavy-tailed return distribution
  • There is a set of sink states $E\subseteq\mathcal{S}$ that terminate every trajectory, which is disjoint from the set of start states.
@@ 3.3 Light-tailed + independence imply $\mathbb{E}[V]\to\infty$
If $U=X+V$ with $X$ and $V$ both light-tailed and $V$ unbounded, and the distribution of U is continuous, and $\pi^{*}(\beta)\triangleq\arg\max_{\pi}\mathbb{E}[U(\pi)]-\beta D_{KL}(\pi,\pi_{0})$, then $\lim_{\beta\to 0^{+}}\mathbb{E}[V(\pi^{*}(\beta))]=\infty$.
3.4 Conditioning as alternate model of optimization
Although we think a KL divergence penalty or cap is the most realistic setting for RLHF, it is not the only model of optimization where heavy-tailedness of the error determines whether catastrophic Goodhart occurs. Consider another model of optimization where $U=X+V$ as before, but we simply condition on $U$ being higher than some threshold $t$ [^1] This could model a satisficing agent that takes random acceptable actions. Then we are interested in the quantity $\lim_{t\to\infty}\mathbb{E}[V|X+V\geq t]$. If we slightly strengthen the heavy-tailedness and light-tailedness assumptions, heavy-tailed error results in catastrophic Goodhart, while light-tailed error results in arbitrarily high expected utility.
3.4.1 Conditioning with heavy-tailed error produces catastrophic Goodhart
Theorem 5.
  • $V$ has a finite mean.
  • The tail of $V$ is sufficiently lighter than the tail of $X$ that $\lim_{t\to\infty}\frac{t^{p}\bar{F}_{V}(t)}{\bar{F}_{X}(t)}=0$ for some $p>1$.
The proof is included in the appendix. It requires expressing the conditional expectation in question as $\frac{\int_{-\infty}^{\infty}vf_{V}(v)\text{Pr}(X>t-v)}{\int_{-\infty}^{\infty}f_{V}(v)\text{Pr}(X>t-v)}$, then partitioning the interval (,)($\infty$,$\infty$)( - ∞ , ∞ ) into four regions and bounding the integrand in the numerator above by a different quantity in each region.
3.4.2 Conditioning with light-tailed error produces arbitrarily high utility
Theorem 6.
Theorem 6 generalizes a consequence of the "Regressional Goodhart Identity" in (Gao et al., , 2023).
4 Experiments
@@
Our theoretical results now raise the question of whether the error in reward models is heavy-tailed or light-tailed in practice.  [^2] Note that distributions over a finite set are bounded and cannot be heavy-tailed in a technical sense, and models with a finite context window have a finite input space. For our purposes, a distribution of reward or error is effectively heavy-tailed if it takes on sufficiently large values and is well-modeled by a heavy-tailed distribution on its support. If we observe the reward distribution to be light-tailed, this is a strong indication that error is light-tailed.  [^3] It is possible for $U$ to be light-tailed while $X$ and $V$ are both heavy-tailed, but this is unusual and we do not expect it to happen in practice.
@@
To empirically test whether the reward is heavy-tailed, we consider two lines of evidence: examining the distributions directly through random sampling and temperature-1 sampling, and finding adversarial token sequences that get high rewards. We examine one small and one medium reward model that performed reasonably well on RewardBench (Lambert et al., , 2023). The small model is an OpenAssistant model based on Pythia 1.4B, and the medium model is Starling 7B-alpha (Zhu et al., , 2023) [^4] https://huggingface.co/berkeley-nest/Starling-RM-7B-alpha.
@@
For random sampling, we sample 30000 length-1024 sequences of uniformly random tokens and observe the distribution of rewards assigned by both Pythia 1.4B and Llama 7B-chat. We also use Llama 7B-chat to generate 16000 length-133 sequences at temperature 1 and observe the distribution of rewards assigned by Starling 7B-alpha.
@@
Because sampling is inefficient at probing the extreme tail, we also find token sequences that optimize Starling 7B-alpha for reward. We considered Greedy Coordinate Gradient (GCG) from (Zou et al., , 2023), a method used to find adversarial suffixes that circumvent jailbreaking, but decided on a faster version of GCG called Accelerated Coordinate Gradient (ACG) from (Haize Labs, , 2024). See Table 1 for ACG hyperparameters.
Parameter Value
Context length 133
Iterations 1000
Candidates per seq. position (k) 3
Annealing starting value 9
Annealing ending value 2
Table 1: Hyperparameters for ACG
@@
4.1 Results
When sampling token sequences, both the Pythia model on random inputs (Figure B.1) and Starling 7B-alpha on Llama-generated inputs (Figure 2) appear approximately normal and, therefore, light-tailed. Starling on random inputs (Figure 1 is ambiguous, with the exponential Q-Q plot having an outlier that could indicate a heavy-tailed distribution, but the Hill estimator is consistent with a light-tailed distribution. Because Llama-7B-chat is a more reasonable base model than a completely random policy, we believe that Starling 7B-alpha is more likely to be light-tailed for the purposes of our theoretical results.
@@ 4.1 Results
When $\alpha$ is small but much larger than $q(x)$, we approximate this to first order as $D_{KL}(P\|Q)\approx\alpha\log\left(\frac{\alpha}{q(x)}\right)$. In Theorems 1 and 3, we prove that when the error is sufficiently heavy-tailed, a policy that gets extremely large reward a small fraction of the time will achieve high expected reward with low KL divergence. This is not the case here because the rewards achieved through ACG were small and the log-probabilities extremely negative. For example, a policy that matches Llama 2-chat’s base reward 99% of the time and uses the highest-reward input generated by ACG $\alpha=$1% of the time will have KL divergence from Llama 2-chat of $\alpha(\log(\alpha)-1339.70)=13.35$ nats, but reward only about $\alpha*(2.2377-0.3329)=0.02571$ greater than the base model, far less than can be obtained with the same KL divergence by conditioning.
@@ 5.1 How likely is catastrophic Goodhart?
  • If we cap KL to a finite value (or dynamically adjust the KL penalty to target a finite KL, as done in Ziegler et al., (2020), then $\mathbb{E}[V]$ is also upper bounded by a finite value (see Theorem 2), and we think it is likely that $\mathbb{E}[V]\approx 0$. Consider a toy model where an AI can adjust three parameters: true quality $V$ of responses, frequency of reward hacking (producing actions with extremely high X), and severity of hacking (value of X on those actions). Adjusting the policy to increase $\mathbb{E}[U]$ without increasing KL increase the severity of hacking while decreasing either frequency of hacking or quality of responses. When $E[U]$ is already large, decreasing quality has much better returns than decreasing frequency. This is similar to Theorems 5, 6 about hard-threshold optimization.
  • Any way we maximize $\mathbb{E}[U(\pi)]-\beta D_{KL}\left(\pi,\pi_{0}\right)$ results in very large values of $\mathbb{E}[U(\pi)]$, and there are a number of arguments that extreme optimization for an imperfect proxy can result in decreased utility due to tradeoffs between $X$ and $V$; e.g., the constrained resource scenario in (Zhuang and Hadfield-Menell, , 2021).